### v1

# Bowen and Wilson Image Class Translation Distance

Abstract
[Abstract] We propose a novel application of image translation networks for image classification and demonstrate its potential as a more interpretable alternative to conventional black box classification networks. We train a network to translate images between possible classes, and then quantify translation distance, i.e. the degree of alteration needed to conform an image to one class or another. These translation distances can then be examined for clusters and trends, and can be fed directly to a simple classifier (e.g. a support vector machine, SVM), providing comparable accuracy compared to a conventional end-to-end convolutional neural network classifier. In addition, visual inspection of translated images can reveal class-specific characteristics and biases in the training sets, such as visual artifacts that are more frequently observed in one class or another. We demonstrate the approach on a toy 2-class scenario, apples versus oranges, and then apply it to two medical imaging tasks: detecting melanoma from photographs of pigmented lesions and classifying 6 cell types in a bone marrow biopsy smear. This novel application of image-to-image networks shows the potential of the technology to go beyond imagining different stylistic changes and to provide greater insight into image classification and medical imaging datasets.
00footnotetext: title footnote…keywords:
CycleGAN, StarGAN, image classification, image translation, interpretable machine learning, dimension reduction1 Introduction
Recent advances in machine learning such as deep convolutional neural networks (CNNs) have brought about revolutionary new capabilities for automated object recognition and classification 1. While the accuracies yielded are impressive, mistakes and misclassifications are still an area of concern, especially in medical applications. CNNs trained on images from one clinic perform poorly on images from another clinic (known as the out-of-domain problem) and trivial changes to an image, e.g., brightness, lighting, or rotation can fool even state-of-the-art classifiers, even when these changes are incorporated as augmentations to the training data 2. Even worse, misclassifications with high confidence scores can mislead even experts into making the wrong call 3. For these reasons, clinicians have been reluctant to make use of AI in routine practice.
One possible solution is to make a better CNN that does not suffer from the out-of-domain problem, a difficult task that has not yet been solved 4. Another solution lies in interpretable/explainable AI (xAI). The idea is if a human can examine a network’s decision-making process they can decide whether to trust the network’s output for a specific case and use that information to make a more informed decision 5, 6. Most of these approaches work with a conventional CNN classifier and analyze hidden layer activations 7. However, even with advances like class activation maps 8, which enable CNNs to highlight regions that contributed to its decision, a CNN still cannot explain what characteristics led to the decision.
Our approach, illustrated in Fig. 1, is enabled by a generative-adversarial 9 method for learning to translate between different two image domains without the need for paired training images 10 or multiple domains 11. Typical applications of image2image are modality transform (I2I-MT) and class transform (I2I-CT). Examples of I2I-MT include style transfer for artistic purposes, such as transforming a photograph into a cartoon-style image 12 and translating between medical imaging modalities, such as rendering a computed tomography (CT) scan from a magnetic resonance image (MRI) for radiation oncology applications 13. In our application, we use the network (in this study, either a CycleGAN or a StarGAN) for I2I-CT and translate each input image into each possible target class. The amount of alteration can be quantified with pixel-wise mean average distance (eg distance), which can be fed to a classifier such as a support vector machine (SVM) 14. Unlike a conventional black-box classifier, a visual inspection of the translated images can draw an observer’s attention to case-specific defining characteristics and potentially expose biases in the training dataset.
A related approach to ours, "Auditing the inference processes of medical-image classifiers by leveraging generative AI and the expertise of physicians" looked to distinguish between melanomas based on patterns from a generative network 15. Their process used an augmented version of a CycleGAN that produced hypothetical images, that a pretrained classifier would then use for predictions. This approach reveals which alterations (e.g. changes in lighting or background skin tone) can change the classifier’s decision, but does not fix any underlying issues with the traditional network. By contrast, our approach explicitly translates an image between possible classes in order to reveal differences between the classes themselves. The amount of alteration to fit each class is then quantified to digest an image into a simple feature vector that can then be fed to a simple classifier, such as logistic regression.
We emphasize that the main focus of this work is not to outperform state-of-the-art classifiers (although the StarGAN experiments offered significant advantages over a conventional end-to-end CNN classifier). Instead, we aim to show how the image2image framework applied to classification tasks can show highlight the network’s interpretation of an image, visualize morphological changes, and uncover learned domain-specific features, without having to analyze hidden layer activations. This in turn can expose both shortcomings of the network architectures and underlying biases in the datasets, as we will show.
We demonstrate the approach first on two class problems with a CycleGAN: a trivial apples versus oranges task, and then the difficult melanoma versus benign pigmented lesion task. We then compared the performance to a conventional CNN and evaluated the shortcomings of that network architecture, moving into tasks with more than two classes. We then discuss overall results including potential training set biases and shortcomings of the underlying CNNs revealed by visual inspection of the results.
2 Methods
We conducted four class transformation (I2I-CT) experiments: 1) 2-class CycleGAN on apples/oranges, 2) 2-class CycleGAN pigmented lesion, 3) 3-class StarGAN on bone marrow cytology, and 4) 6-class StarGAN on bone marrow cytology. For each, the performance of the image-to-image approach was compared with a baseline CNN end-to-end classifier (EfficientNet 16, pre-trained on ImageNet 17). For the 2-class problems, to measure performance while accounting for class imbalance we used the area under the receiver operator curve (AUROC) 18. For the multi-class problems, we looked at traditional overall accuracy and then individual class accuracy in a confusion matrix. The resulting generated images were also visually compared with the source images.
2.1 Classification Datasets
2.1.1 2-class problems
The apples/oranges dataset is a subset of the larger ImageNet dataset 19 compiled by the CycleGAN authors 10, from which we randomly selected a training set of 995 apple and 1019 orange images; our test set had 266 apple and 248 orange images.
The pigmented lesions dataset consisted of dermoscopic photographs from the International Skin Imaging Collaboration (ISIC) 2020 challenge dataset 20. Images were collected from the Hospital Clínic de Barcelona, the Medical University of Vienna, Memorial Sloan Kettering Cancer Center, Melanoma Institute Australia, the University of Queensland, and the University of Athens Medical School. There were 30,928 benign images and 584 malignant images. To mitigate the class imbalance problem 21, we used a random subset of 2,584 images with an 80/20 train/test split (453 malignant and 1614 benign images in the training set and 131 malignant and 386 benign images in the test set). The baseline CNN and CycleGAN networks used the same data splits for performance comparisons.
While initial training used the ISIC data from 2020, out-of-distribution performance was tested with malignant images from the 2016 and 2017 ISIC challenges 22, 23, which is comprised of an entirely separate set of images, and additional benign images from the ISIC 2020 challenge that were excluded from our initial subset 20.
2.1.2 Multi-class
For the multi-class experiments, we used the Bone Marrow Cytology in Hematologic Malignancies Data 24. This had 21 different classes. For the 3-class model, we selected three categories: blasts, erythroblasts, and lymphocytes. These classes were selected because of the large availability of images. The number of training images was set at 5,000 images initially for each class (randomly selected), to eliminate the class imbalance problem during training. This ensured that each class was equally likely to be seen during training. The remaining images of each class were used for testing. Training consisted of 15,000 images, 5,000 of each respective class. 50,610 images were used for testing (Blasts: 6,973, Erythroblasts: 22,395, Lymphocytes: 21,242).
For the 6-class model, we used those three classes and added segmented neutrophil, plasma cell, and promyelocyte. These additional three classes were selected because they met the requirements of having more than 5,000 total images. 5,000 images from each of these classes were added to the images of the previous three classes, and the remaining images added to the test set. (Segmented Neutrophil: 24,434, Plasma Cells: 2629, Promylelocyte: 11,994). To test robustness, we wanted to see the capabilities on a larger dataset compared to the 3-class example.
An additional experiment leveraged the entire 6-class dataset with an 80/20 train test split using a weighted sampler to mitigate the class imbalance problem. We used PyTorch’s WeightedRandomSampler during the data processing phase to calculate and create class weights. The inverse frequency of each class was given as the weight for each image based on its target. , where is the class frequency.
2.2 Baseline CNN classifier
For each dataset, a baseline CNN was implemented using transfer learning. An EffecientNetB3 16 was pre-trained on ImageNet 17. Training augmentations included vertical and horizontal flip, random brightness and contrast, optical distortion, grid distortion, elastic transform, contrast limited adaptive histogram equalization, cutout, and color jitter, using the Albumentation package 25.
To adjust for class imbalance, class weights were used with the datasets with a traditional 80/20 split. This was calculated as . For the datasets with already an equal number of images for each class during, the CNN used the same dataset for comparison purposes with the StarGAN.
The CNN classifiers were trained for fewer iterations than the StarGAN networks due to a CNNs being prone to overtraining. The network used an early stopping callback to monitor the validation set loss with a patience of 200 epochs. The CNN was trained to be a baseline, achieving results that could be comparable to the StarGAN.
2.3 Image Translation Networks
2.3.1 2-class: CycleGAN network
The CycleGAN model was a TensorFlow implementation 26 of the original PyTorch CycleGAN 10 for 256x256 images. The only modification we applied was the addition of dropout to the residual layers.
Initially, we used the default values of and . Both the generators and discriminators used the optimizer Adam with a learning rate of and . Models were trained for 60,600 iterations.
2.3.2 Multi-class: StarGAN network
The StarGAN model was the first version of the StarGAN 27. Multi-domain translations are possible with a single network. Labels are given to the different domains and a single generator and discriminator are used to learn the mappings for each image into every domain.
The StarGAN did not originally use an identity loss or calculate translation distances. We implemented a version of identity loss as seen in the CycleGAN 10 and tuned that to improve translation distances and classification. The Bone Cytology and Pigmented lesion experiments used a . (Note: this is smaller than in the CycleGAN because of the difference between the two codebases in using mean absolute error versus sum absolute error in their loss functions). The model was trained for a minimum of 300,000 iterations for the experiments. Increasing the number of classes and samples during training and testing, required more iterations to achieve comparable image quality. A batch size of 16, an image size of 128, and the rest of the default parameters the same as the original StarGAN.
2.4 Image translation distance metrics
To quantify image changes and generate the translation distance feature set, an input image is fed through all generators, producing a set of class-specific hypothetical images, . Using the input images and the generated images, we calculated translation distance (pixel-wise mean average difference) for each class, i.e. , to quantify how much an image needed to change to fit into each class. We expect images to have small translation distances for classes to which they belong and larger translation distances for classes to which they do not belong.
For two classes, the translation distances were visualized with scatterplots, and a translation ratio was calculated, , as a simple estimate of the probability of an image belonging to class . Finally, simple classifier was trained to separate images based on their translation distance coordinates . Most of the experiments used a linear support vector machine (SVM) 14 (with category weighting when dealing with the the pigmented lesions and full bone marrow cytology datasets), but in the 6-class experiment we compare several other classifiers.
The predictions from the CycleGAN+SVM were used to plot the receiver operating characteristic (ROC) and to calculate AUROC as a measurement of the CycleGAN’s ability to separate images by class. As described later, we sought to maximize this CycleGAN+SVM AUROC by tuning the CycleGAN identity loss and cycle consistency loss .
The predictions from the StarGAN+SVM were used to create confusion matrices to monitor class accuracy, specificity, and sensitivity and tune .
3 2-class Results
3.1 Apples and Oranges
The separability of apples and oranges in translation distance space can be seen in the scatterplot (Fig. 2(b)) and translation ratio histogram (Fig. 2(c)). The traditional CNN classified apples and oranges with an AUROC of .9922 and .9825 on the training and test sets, respectively. The CycleGAN+SVM achieved an AUROC of .9926 and .9545 on the test and training sets, respectively. By changing to , the AUROC increased to .9971 and .9705 on the training and test sets. Overall, the traditional CNN only outperforms the CycleGAN+SVM approach by a marginal 1.2%, as measured by the test set AUROC, demonstrating the potential of using image translation networks for classification.
Fig. 3 shows a selection of images produced from the apples (class A) to oranges (class B) CyleGAN model with , . The top row of Fig. 3 shows images of reasonable translation quality that were correctly classified by the CycleGAN+SVM. The middle row shows images that failed to translate morphology in the opposing class but translation distances still led to correct classification. The bottom row shows uncharacteristic source images that led to translation or classification failures. Visual inspection shows features such as fruit color, background color, and texture that were changed or added to the original image to produce hypothetical images for each class.
3.2 Pigmented Lesions
Classifying melanoma versus benign pigmented lesions is relatively more difficult than apples versus oranges. First, we performed hyperparameter tuning to maximize the separation of classes (with an SVM weight of 3 to adjust for the class imbalance). After evaluation models with different combinations of and , we found that and were the best parameters and that identity loss is necessary for preventing the network from changing an image if it is already in the correct domain. (With this class translation task, cycle consistency loss turned out to prevent any meaningful alterations to the images). The CycleGAN+SVM achieved a .8044 AUROC, compared to .8536 AUROC achieved by the CNN classifier. Both approaches are within the range of AUROC reported on the melanoma versus benign dermoscopy task, both for dermatologist visual inspection, where AUROC ranges from 0.671 28 to 0.79 29, with high levels of sensitivity and specificity (70%s and higher), and for CNNs, where AUROC from state-of-the-art networks ranges from 0.654 to 0.949 4 depending on the dataset.
The translation distance scatterplots (a) and (b) in Fig. 4 show that the separation between the malignant and benign pigmented lesions is not as pronounced as in apples versus oranges (Fig. 2). Three regions in translation distance space can be identified: (1) contains benign lesions that are well-separated from the majority of melanomas; (2) contains melanomas that are well-separated from the majority of benign lesions; and (3) contains overlap between benign and malignant lesions. Upon further examination, testing with a larger number of benign images and malignant images from multiple years of the ISIC database (Fig. 5), we found that region (1) consists mostly of images labeled "Single Image Expert Benign" as opposed to "Histopathology Benign" – in other words, these images were considered unambiguously benign by a dermatologist, not requiring a biopsy to confirm. Furthermore, we find that region (1) tends to exclude melanomas, even with a test set expanded to include images from other years of the ISIC database.
For visual inspection, we first show a selection of images (Fig. 6) that were correctly classified by the CycleGAN+SVM at a decision threshold of 0.5. In these cases, rendered images from the correct class tend to look more like the source image in terms of the pigmented lesion’s color (brown/black), texture, and morphology. However, the generators also tend to alter other characteristics such as background skin color (in these cases, pale versus pink), scalebar markings, and vignetting (darkening around the edges due to the imaging optics or lighting conditions). From this, it would appear that the generators have learned to prefer rendering benign lesions with a pink background skin tone, and melanomas with a pale skin tone, scalebar markings, and vignetting. This suggests the presence of these artifacts may be biased towards one class or another in the training data, and may be a source of error in training conventional classifiers.
Secondly, a selection of images misclassified by the traditional CNN is shown in Fig. 9. In these cases, though the CycleGAN+SVM does not always arrive at the correct classification, the generated images provide insight into artifacts and possible causes for misclassification. For example, the benign images in Fig. 9 show the same tendency of the network to consider benign images as having a more pink background tone and melanomas as having lighter skin tone and vignetting. As discussed above, this may reflect biases in the training dataset (further examples can be seen in the supplementary materials). For the malignant images in Fig. 9 (b), while the traditional CNN gave a low probability of melanoma, the generated malignant images appear visually much closer to the source images input malignant lesions, as is also reflected in the translation distances.
Thirdly, a selection of images misclassified by the CycleGAN+SVM is shown in Fig. 9. In these cases, a visual inspection affords more room for interpretation and making an informed decision than would a black-box classifier. Consistent with earlier observations, it appears that benign lesions (Fig. 9 (a)) with pale skin background, scalebar markings, and vignetting tend to be misclassified as melanomas. In the case of malignant lesions (Fig. 9 (b)), though the SVM incorrectly labeled these benign, the source images are more visually similar to the malignant renderings.
Finally, a selection of generator failures/hallucinations are shown in Fig. 9, in which rendered images in one domain or another show severe deviations from the source image. However, in each of these cases, both the CycleGAN+SVM and visual inspection lead to correct classification. Note again the tendency towards pink skin tone in the rendered benign images and pale skin and vignetting in the rendered malignant images.
4 3-class Results
4.1 Bone Marrow Smear Cytology
To accommodate more than two classes, we switched from a CycleGAN to a StarGAN.
Utilizing the Bone Marrow Cytology in Hematologic Malignancies dataset, we achieved 95% training accuracy, and a 92% test accuracy in a 3-class identification task, predicting classes using the smallest translation distance. This showcases the method’s potential on larger tasks. This was a marginal improvement over the traditional CNN, which achieved an 89% test accuracy.
To get a better breakdown of class performance we evaluated the confusion matrices of both networks(Fig. 7 Although, the overall accuracy was close, there were differences in the class accuracies. Erythroblast class accuracy was similar (95% EfficientNet vs 96% StarGAN) with the StarGAN having a slight edge. The StarGAN also had a 3% higher accuracy for the lymphocytes and also was able to improve blast accuracy by 10% over the EfficientNet.
Blasts and lymphocytes are more likely to be mistaken for each other by both models. The EfficientNet had more blast false negatives and lymphocyte false positives than the StarGAN. Although, the StarGAN decreased the amount of false positives and negatives, the same underlying trend remains that blasts and lymphocytes are more likely to be mistaken for each other compared to erythroblasts.
In image samples in Fig. 8, we can see some morphology changes that occurred for each of the images to fit into each class. The top row shows an image of a blast, that was labeled correctly by the image2image network, but misclassified by the CNN. The original blast is darkly stained, and it is difficult to distinguish the nucleus from the cytoplasm. The translation to a blast (in-class) retains these characteristics. The translation to an erythoblast resulted in a centered round nucleaus surrounded by a lighter-stained cytoplasm, while retaining the overall cell. Lastly, the translation to a lymphocyte changed the hue to a darker purple.
The middle row of Fig. 8 shows an erythroblast that was correctly identified by both networks. Translation to an erythrocyte (in-class) retains the shape of cell and nucleus, and the contrast between cytoplasm and nucleus. Translation to a blast resulted in a loss distinction between the nucleus and cytoplasm, similar to the blast in the top row. Translation to a lymphocyte expanded the nucleus.
The bottom row of final image panel shows a lymphocyte with a lighter-staining cytoplasm than the previous two examples. Translation to a lymphocyte (in-class) retains these characteristics. Translation to a blast image enlarged the nucleus and darkened the cytoplasm, consistent with the blast shown in the top row. Translation to an erythroblast, on the other hand shrank the nucleus and maintaned a lighter cytoplasm.
In these examples, the StarGAN followed by a translation distance classifier was able to correctly classify the images. The translation distances mirror showcases the classification decision, with the smallest distance being the classification. The CNN was able to classify erythroblast and lymphocyte correctly, but was unable to classify the blast.
We also highlight some image translation failure cases. Fig. 9 shows a correctly classified cell, however, there are hallucinations in the produced images, especially in the translation to blast. These hallucinations alter the image in such a way that the image no longer looks realistic. This is important in showing where the network might be overgeneralizing or is undertrained, but we note that the in-class translation still was achieved with lower translation distance than the other classes.
5 6-class Results
5.1 Bone Cytology
Expanding to six classes, the image2image network was able to achieve 90% testing accuracy (91.96% with a custom logistic regression classifier on translation distance) by classifying the image using the smallest translation distance. The traditional end-to-end CNN, by comparison achieved only 86% accuracy. (Fig. 1) Individual class accuracy ranged from 86%-97%.(Fig. 9)
The concern from the 2-class and 3-class problems was the potential that the network would not be able to generalize well or be expanded to more classes. However, this experiment showed that accuracy was still able to remain high with an increased number of classes and samples.
The original three cytology classes were more morphologically similar: blasts, erythroblasts, and lymphocytes all have large, darkly stained nuclei. Lymphocytes differ by having a thin rim of lighter-stained cytoplasm. The three added classes added more variety to the morphology in the available images. For instance, the samples of the segmented neutrophil tended to have a characteristic clear cell body (chromophobic cytoplasm) and segmented nuclei. As a result, the generated images for that class needed to either add or modify features to resemble that characteristic. In Fig. 10, the segmented neutrophil has more morphological changes than the other images. As a result, the translation distance is greater and it does not closely resemble the input image.
Despite high overall accuracy, and high class accuracy, the network still made some classification errors. In Fig.11, the network incorrectly classified a blast as a promelocyte. This was an error that the traditional CNN made as well. The end-to-end CNN in this case has a misleading confidence of 99.8% that the image is a promyleocyte. This high-confidence prediction makes it hard to question the decision, but in this case, it is an overconfident and incorrect decision. The promyelocyte had the lowest translation distance and was selected as the class. However, a quick visual inspection shows that the hypothetical blast and promyelocyte both look visually similar to the input blast image, and both have small translation distances, indicating correctly, a much lower confidence than the end-to-end CNN suggests.
For the 6-class experiment, the translation from the identity mapping can be better visualized. This principle on identity loss in CycleGANs 10, emphasizes that an image should not need to adjusted if it stays within its own class. In Fig. 9, showcases images that were correctly classified. For the segmented neutrophil image, the identity mapping preserves the characteristics of the original images, whereas the other classes modify the image. Additionally, even where the image is incorrectly classified using translation distance alone in Fig. 9, images translated into their own domains preserve original characteristics. In this panel, other images do not modify the image either, which leads to closer overall translation distances between classes. Classes that did show alterations had much higher translations distances and were less likely to be considered as classifications initially.
Finally, we note that conducting the same experiment with PyTorch’s Random Weighted Sampler to solve the class imbalance problem and achieves comparable accuracies to these experiments.
5.2 Translation Distance Classifiers
In the original experiments with a CycleGAN, we applied an SVM for the final classification. Adjusting the SVM, we were able to achieve high accuracy. However, in the multi-class classification, the translation distance was better able to reflect class differences. As a result, although an SVM could improve classification results, translation distance alone could be used for the final decision.
| Translation Distance Classifier | Training Accuracy (%) | Testing Accuracy (%) |
| Custom Linear Regression | 91.92 | 91.96 |
| Scikit Logistic Regression | 91.87 | 91.90 |
| Scikit SVM Linear Kernel | 91.40 | 91.71 |
| Scikit SVM RBF Kernel | 91.42 | 91.65 |
| Scikit SVM Poly Kernel | 90.52 | 91.29 |
| Lowest Translation Distance | 91.03 | 90.07 |
| Multilayer Perceptron | 88.37 | 88.72 |
| (EfficientNetB3 end-to-end) | 70.27 | 86.07 |
To demonstrate and determine the best accuracy, we tested a variety of simple classifiers. The translation distance classifiers all had similar accuracies, and their confusion matrices were very similar with slight differences in individual class accuracies. Translation distance alone does not use another network; instead, the smallest distance is selected as the predicted class. Adding a classifier can improve accuracy with a slight cost of interpretability allowing the networks to maximize accuracy. A custom logistic regression model with a softmax activation function is the most effective for the 6-class problem, however, most of the classifiers shown in Tab. 1 can achieve acceptable accuracy results.
Although overall accuracy was similar between the classifiers, there were some differences between class accuracies, both within the networks and compared the other classifiers. In Fig. 9), blasts have the lowest class accuracy at 86% and segmented neutrophils have the highest with 97%. Selecting different auxilary classifiers, results in slightly different class accuracies and ranges.
In addition to outperforming the overall accuracy of the traditional network, the custom logistic regression classifer resulted in higher individual class accuracies. The traditional network achieved only a 65% blast accuracy (Fig. 9), but the custom logistic classifier returned 65% (Fig.9). Similiarily, the lymphocyte had 71% and 87% accuracies, respectively.
6 Conclusion
Using image-to-image networks for classification is built on the same idea behind the identity loss in CycleGANs 10–an image translated into its own domain/class should not require any changes. Conversely, images translated into domains/classes to which they do not belong require more changes. Thus by comparing visually the source and generated images, or quantifying image translation distances (e.g. ) it is possible to determine the class of a source image . This method reduces each input image into a point in translation distance space, , which can be analyzed for clusters and trends. Remarkably, in the two class experiment with melanoma, without having been explicitly trained to do so, the network described here grouped pigmented lesions in a similar manner to a dermatologist’s call on whether to perform a biopsy (see Fig. 5 and discussion). Furthermore, the overlap between benign lesions that had to be biopsied and melanomas raises the question of whether it is even possible to distinguish some types of benign lesions from melanomas. While the drive for improved accuracy often assumes all benign lesions are distinguishable from melanomas (and vice-versa) from a dermoscopy photograph alone, this is just not the case, as documented by earlier investigations in the limitations of dermoscopy 30. We note that other networks and training approaches can in principle be used, and may provide better results, e.g. contrastive unpaired translation 31, DRIT 32, cyclic autoencoders 33, transformers 34, and stable diffusion models 35.
Unlike existing image dimension reduction methods, e.g. principal component analysis 36, 37, convolutional autoencoders 38, stochastic neighbor embedding 39, etc., the dimensions of this new translation distance space are straightforward and interpretable as the amount of change an image must undergo to fit into a given class. The translation distances are independently significant and can be used as features in downstream classifiers, whose inputs and weights are then much easier to interpret than end-to-end classifiers.
The image2image framework was able to achieve comparable accuracies, even outperforming a traditional CNN. Furthermore, it provided greater insights into how the network was making decisions. In the bone cytology experiment, classification and visual errors could be more easily inspected. The baseline CNN classifier, an EfficientNetB3, is one that resulted from extensive architecture hyperparameter tuning, and also made use of transfer learning. The image translation networks, CycleGAN and StarGAN, had none of those advantages, relying on small identity loss adjustments. This suggests the possibility that with further development, image-to-image networks may achieve better accuracy on classification tasks.
The network struggled to perform without an adjustment for class imbalance though, as some groups were over represented. Selecting an equal number of samples from each class during training solves this problem. However, it does limit the number of images seen. To increase the samples during training, weighted sampling allows the traditional 80/20 split to be used. The class distribution problem can be solved either way.
The translation networks are trained on the task of creating realistic images that can deceive the discriminator. Using the translation distance to classify images, adds to the functionality of the already robust discriminator within the image2image network. The training and testing results of the classification aspects, are secondary metrics of an already trained network. Consequently, the network retains comparable training and testing accuracies even when trained on a reduced dataset was selected to mitigate the class-imbalance problem. CNNs have been observed struggle to apply their performance outside the training set. Therefore, the StarGAN’s ability to achieve comparable accuracy in a larger test set is promising.
With this approach, translation distances can be used for classification or can be analyzed with a simple classifier (e.g. an SVM, a linear classifier or any number of neural network approaches can be used). The resulting generated images can be visually inspected to drive more informed decisions and uncover insights into the training dataset. For example, consider the artifact of vignetting (darkening at the edge of images). A human observer would overlook this as irrelevant to detecting melanoma, but a training dataset that exhibits certain artifacts more frequently in one class or another would easily mislead a conventional classifier. An image-to-image approach has the potential to reveal the severity of artifacts on a case-by-case basis: an artifact’s presence in the source and both generated images indicates a robust translation result, whereas an artifact’s absence in one of the generated images indicates it was mistaken for a defining class-specific feature. Such mistakes may result from having artifacts that are unequally distributed between training set classes, as we uncovered with vignetting, skin tone, and scalebar markings in the melanoma versus benign dataset. Such insights from applying image-to-image networks to difficult classification tasks will hopefully inform efforts to produce better training datasets and in turn, improve outcomes with conventional classification CNNs.
Finally, the image2image framework can help bridge the interpretability gap between technologies such as (AI, and microscopy techniques) with their users, clinicians. While the clinician-in-the-training-loop approach yields impressive improvements in accuracy and robustness against artifacts 15, our approach is less labor intensive and provides clinicians an opportunity to examine visual differences between classes at inference (point of care). Both approaches are complementary and add value to AI processing and evaluation of medical imaging by working alongside, rather than to replace clinicians, and have potential for both improving diagnoses and in training human observers to recognize subtle differences between classes. Furthermore, the use of image2image for classification can reveal biases, morphological similarities and differences and show new patterns in the data, creating new lines of inquiry.
7 Acknowledgments
We thank Priya Wolff (University of Colorado School of Medicine) and Yevgeniy Seminov (Mass. General Brigham) for helpful conversations and suggestions, and Bill Carpenter (Colorado State University) for technical support with computational resources. Experiments were carried out on the Riviera cluster (Data Science Research Institute, Colorado State University). This work was funded by the Boettcher Foundation through a Boettcher Collaboration Grant and a Boettcher Educational Enrichment Grant.